how to do this problem?
the perimeter of a rectangle is 28 cm. one side is 10 cm longer than the other side. find the lengths of the sides.

Answer by TimothyLamb(4379)

2L + 2w = 28
L = w + 10
---
2L + 2w = 28
2(w + 10) + 2w = 28
2w + 20 + 2w = 28
4w = 8
---
w = 2
L = 12
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php